OpenStudy (anonymous):

The incidence of cystic fibrosis, a recessive genetic disorder in the Caucasian population of United States, is 1 in every 2,500 individuals. Find the number of heterozygous carriers. (p + q = 1, p2 + 2pq + q2 = 1)

OpenStudy (matt101):

Since cystic fibrosis is a recessive disorder, 1/2500 = p^2, so p = 1/50. Since p + q = 1, q = 1 - (1/50) = 49/50. The number of heterozygous carriers is 2pq = 2(1/50)(49/50) = 98/2500, or 98 in every 2500 people.